# Scalable Multitask Learning Using Gradient-based Estimation of Task Affinity

### v1

@@ 3.2. Gradient-based Estimation
We now describe our algorithm, which builds on the above linearization property, by using logistic regression with gradients as features. It also includes dimension reduction, as described below.
@@
The main idea is to solve a logistic regression problem with $g_{i}$ being the feature vector and $y_{i}$ being the response label. However, keep in mind that the dimension of $g_{i}$ is the same as the number of parameters in a neural network, which could be tens of millions. Thus, we introduce a dimension reduction procedure that does not lose much precision.
@@
Lastly, we set $\hat{W}_{S}$ as $P\hat{W}_{d}+\theta^{\star}$ to map the projected solution back to the $p$-dimensional space. $\hat{W}_{S}$ is the estimated model parameter for fine-tuning $\theta^{\star}$ with task subset $S$.
@@
Typically, $M=O(1)$, while $m=\Omega(n)$ or even $O(n^{2})$ in downstream use cases. Thus, our estimation algorithm reduces $\Omega(n)$ full-model training to only $O(1)$. The tradeoff is that we require $O(n)$ gradient evaluations (to retrieve the gradients on all tasks) plus solving logistic regression $m$ times. As we will show below, the random projection helps reduce the dimension of the logistic regression problem to $O(\log p)$ dimension, which is much cheaper. This is in terms of the asymptotic complexity. In Section 5.2, we materialize the constants to compare the number of FLOPs during training.
@@
We note that the above procedure can be extended to deal with multiple classifications. This requires setting up one prediction vector for each class; The rest remains the same. The procedure also applies to regression by using mean squared error instead.
@@ 3.3. Error Bounds
We now show that the error introduced by approximations in Grad-TAE is bounded. Specifically, we use the Johnson-Lindenstrauss Lemma to argue that as $d$ increases, the random projection yields a minimizer whose quality is not much worse than the solution without the projection. We will assume that the averaged Taylor’s expansion error is at most $\delta$ across the entire data set of every task. Additionally, we assume that the search procedure occurs within a bounded space of radius $D$. Lastly, in the pretrained initialization, each gradient vector’s Euclidean norm is at most $G$. With these conditions, we state the error bounds for Grad-TAE as follows.
@@
The proof, given in Appendix A, uses the Johnson-Lindenstrauss Lemma (johnson1984extensions). In particular, using the fact that the logistic loss is $1$-Lipschitz continuous, we can relate $\hat{L}(\hat{W}_{S})$ to $\min\hat{L}(W)$. The errors introduced by random projection and Taylor’s expansion can be bounded using the JL Lemma and the bound on Taylor’s expansion error, respectively. Further, our experiments in Table 1 suggest that $\delta$ is relatively small in practice. Thus, as $\epsilon$ goes to zero, Eq. (6) guarantees the gap between $\hat{L}(\hat{W}_{S})$ and $\min\hat{L}(W)$ will be small.
@@ 4. Task Affinity based Grouping
We now describe a clustering algorithm to partition the $n$ tasks into $k$ disjoint subsets. Given an $n$ by $n$ task affinity matrix $T$, we will find a clustering that maximizes the average density of all clusters. Concretely, let $C_{1},\dots,C_{k}$ be a disjoint partition of $[n]$. Let $v_{1},\dots,v_{k}$ be a 00-$1$ vector indicating whether a task is in one cluster or not. The average density of this clustering can be written as:
@@
We design a Semi-Definite Programming (SDP) relaxation and then round the SDP solution to a clustering. Let us denote the assignment variables as an $n\times k$ matrix $V$, such that each entry $V_{i,j}$ indicates whether a task $i$ belongs to a cluster $j$, for every $i=1,\dots,n$, $j=1,\dots,k$. Moreover, let the $j$th column of $V$, which is the characteristic vector of the $j$-th cluster, be denoted as $v_{j}$. Under this assignment, the sum of $V_{i,j}$ across any task $i$ must be one, as we allow one task to be assigned in a single group. By contrast, the sum of $V_{i,j}$ across $C_{j}$ is the number of tasks assigned to $C_{j}$, which is at least one.
@@
A natural question is using alternative algorithms such as spectral clustering or Lloyd’s clustering. We find that these algorithms are not as robust as the SDP relaxation because the scale of the loss values varies across rows for different tasks. We describe a toy example to illustrate. Suppose $T$ is a $6$ by $6$ matrix involving three clusters $C_{1},C_{2},C_{3}$ of size $2$ each. The affinity in $C_{1}$ is $7$, while the affinity scores in $C_{2}$ and $C_{3}$ are $20,19$, respectively. We find that both spectral clustering and Lloyd’s clustering will group $C_{2}$ and $C_{3}$ together, while the SDP relaxation manages to separate them apart. See Figure 2 for an illustration. For this reason, we use the SDP relaxation in Grad-TAG.
@@
A natural question is whether one can quantify the approximation ratio of the SDP relaxation (10). Although this is a well-studied problem in approximation algorithms (ailon2008aggregating), task affinity violates the metric condition typically required in order to obtain guarantees in this literature. In particular, the triangle inequality $T_{i,j}+T_{j,k}\geq T_{i,k}$ is violated. It is possible that by making an assumption regarding intra-cluster separation (see, e.g., awasthi2015relax), one might be able to analyze the SDP theoretically. This is left for future work.
@@ 5. Experiments
Our experiments show that Grad-TAE approximates the true task affinities (based on full model training) within a relative error of less than 2.7%, while using less than 3% of the computational cost of full training. Further, Grad-TAG achieves comparable downstream accuracy to existing methods in two canonical applications, multi-label classification on graphs and language model fine-tuning, while using 32.8$\times$ fewer FLOPs. Lastly, we discuss the parameters and the steps as part of our algorithm, including the comparison with alternative clustering.
@@ 5.1. Experimental Setup
We note that our algorithm applies to a wide range of multitask learning scenarios. For a representative evaluation, we focus on multi-label prediction on graphs, and language model fine-tuning. In the first setting, each labeling task corresponds to a subgraph within a graph. Given a seed set of each labeling as the training set, the goal is to identify the remaining nodes of the subgraph. This can be cast as multitask learning, by viewing each labeling as a binary classification task. The objective is to optimize the average accuracy of all the labeling tasks.
@@
For fine-tuning language models, we use two text classification datasets from SuperGLUE (wang2019superglue), specifically RTE and WiC. Each dataset includes 100 instructions, with 10 sourced from bach2022promptsource and 90 generated using the automatic instruction generation method in (zhou2022large). Thus, each dataset has 100 tasks in total, each corresponding to fine-tuning with one instruction. We use T5-Base (roberts2019exploring) as the encoder for the MTL model. The choice of this encoder is without loss of generality, as we expect similar results to hold on other encoders.
@@
We assess the accuracy of estimated task affinity by measuring the distance between our estimated task affinities and the task affinities computed from fully trained models.
@@ 5.2. Task Affinity Estimation
We now report the results from running our estimation procedure. We regard the task affinity scores computed from fully trained models as the target, denoted as ${T}^{\star}$. Then, after running Grad-TAE, we compute the affinity matrix $T$, and measure the relative distance between $T$ and ${T}^{\star}$ as:
@@
To explain our findings, we set the number of meta-initializations to $M=1$ and vary the projection dimension $d$ among 50, 100, 200, and 400. We note that all these values yield an estimation of $T^{\star}$ within 11% distance. As expected, increasing $d$ leads to better estimation. After $d$ increases above $200$, the distance metric also stabilizes to around 5.7%. Thus, we set $d$ as 200 in the remaining experiments. As a remark, this is approximately $15\log(p)$, where $p=683,370$ in this experiment, aligning with our analysis in Proposition 3.3. Remarkably, under this setting, Grad-TAE uses 3.5 GPU hours and achieves 130$\times$ less computation compared to fully-trained models!
Next, we fix $d=200$ while increasing $M$ up to $9$. This further reduces the distance metric to 5.4%, with 45.0$\times$ less compute cost. We observe diminishing returns from ensembling, once $M$ goes beyond $5$. Thus, we will set $M$ as $5$ in the remaining experiments. This uses 17.6 GPU hours and 44.9$\times$ less computation than fully-trained models.
@@
We have shown that Grad-TAE significantly reduces the computational cost in task affinity estimation. To verify that these efficiency gains are consistent across different data modalities, we apply Grad-TAE to a text classification dataset (RTE) and an image classification dataset (DomainNet) (peng2019moment). The RTE data set contains 100 tasks. We use T5-Base and compute higher-order task affinity with 2000 subsets of size 10. The DomainNet data set contains 6 tasks. We use ResNet-50 and compute higher-order task affinity with 20 subsets of size 3. On the two data sets, our algorithm reduces computation by 42.6$\times$ and 9.5$\times$, respectively, compared to computing true higher-order task affinities, while incurring less than 3% relative error. The smaller speedup in the image dataset is due to the fewer total models trained on task subsets.
@@ 5.3. Comparison for Task Grouping
We set up a wide range of baselines covering heuristic solutions and recent optimization techniques.
@@ 5.3.3. Fine-tuning language models
Next, we report the results from fine-tuning language models (T5 base) on text classification with $n=100$ instructions. We again use our algorithm to estimate higher-order task affinity scores and apply SDP clustering to group tasks. We illustrate our results in Figure 4(b) while deferring the complete comparison to Appendix C. We use $1-$ accuracy as the error rate on the text classification datasets. In particular, our algorithm outperforms single-task learning by 1.9%.
@@ 6. Conclusion
This paper designs an efficient estimation algorithm to compute task affinity scores. The main idea is to first pretrain a meta-initialization on all tasks and then use the initialization’s gradients to estimate the fine-tuned model parameters for a particular task combination using logistic regression. A random projection is applied to the gradients to reduce the dimension of the regression. Then, we design a robust clustering algorithm to accompany the task affinity estimation, which together yields an efficient multitask learning algorithm. Experiments show that the algorithm can scale to as many as $500$ tasks on very large graphs while accurately approximating the true task affinity scores. The overall algorithm gives the best tradeoff between computation and performance compared to existing multitask learning methods.
@@ Acknowledgement
Thanks to the anonymous referees for their comments. This research is supported in part by NSF award IIS-2412008 and Northeastern University’s Transforming Interdisciplinary Experiential Research (TIER) 1: Seed Grant/Proof of Concept Program.

### v2

@@ 3.2. Gradient-based Estimation
We now describe our algorithm, which builds on the above linearization property, using logistic regression with gradients as features. It also includes dimension reduction, as described below.
@@
The main idea is to solve a logistic regression problem with $g_{i}$ being the feature vector and $y_{i}$ being the response label. However, recall that the dimension of $g_{i}$ is the same as the number of parameters in a neural network, which could be tens of millions. Thus, we introduce a dimension reduction procedure that does not lose much precision.
@@
Lastly, we set $\hat{W}_{S}$ as $P\hat{W}_{d}+\theta^{\star}$ to map the projected solution to the $p$-dimensional space. $\hat{W}_{S}$ is the estimated model parameter for fine-tuning $\theta^{\star}$ with task subset $S$.
@@
Typically, $M=O(1)$, while $m=\Omega(n)$ or even $O(n^{2})$ in downstream use cases. Thus, our estimation algorithm reduces $\Omega(n)$ full-model training to only $O(1)$. The tradeoff is that we require $O(n)$ gradient evaluations (to retrieve the gradients on all tasks) plus solving logistic regression $m$ times. As shown below, the random projection helps reduce the dimension of the logistic regression problem to $O(\log p)$ dimension, which is much cheaper. This is in terms of the asymptotic complexity. In Section 5.2, we materialize the constants to compare the number of FLOPs during training.
@@
The above procedure can be extended to deal with multiple classifications. This requires setting up one prediction vector for each class; The rest remains the same. The procedure also applies to regression by using mean squared error instead.
@@ 3.3. Error Bounds
We now show that the error introduced by approximations in Grad-TAE is bounded. Specifically, we use the Johnson-Lindenstrauss Lemma to argue that as $d$ increases, the random projection yields a minimizer whose quality is not much worse than the solution without the projection. We will assume that the averaged Taylor’s expansion error is at most $\delta$ across the entire data set of every task. Additionally, we assume the search procedure occurs within a bounded space of radius $D$. Lastly, in the pretrained initialization, each gradient vector’s Euclidean norm is at most $G$. With these conditions, we state the error bounds for Grad-TAE as follows.
@@
The proof, given in Appendix A, uses the Johnson-Lindenstrauss Lemma (johnson1984extensions). In particular, since the logistic loss is $1$-Lipschitz continuous, we can relate $\hat{L}(\hat{W}_{S})$ to $\min\hat{L}(W)$. The errors introduced by random projection and Taylor’s expansion can be bounded using the JL Lemma and the bound on Taylor’s expansion error, respectively. Further, our experiments in Table 1 suggest that $\delta$ is relatively small in practice. Thus, as $\epsilon$ goes to zero, Eq. (6) guarantees the gap between $\hat{L}(\hat{W}_{S})$ and $\min\hat{L}(W)$ will be small.
@@ 4. Task Affinity based Grouping
We now describe a clustering algorithm to partition the $n$ tasks into $k$ disjoint subsets. Given an $n$ by $n$ task affinity matrix $T$, we will find a clustering that maximizes the average density of all clusters. Concretely, let $C_{1},\dots,C_{k}$ be a disjoint partition of $[n]$. Let $v_{1},\dots,v_{k}$ be a 00-$1$ vector indicating whether a task is in one cluster. The average density of this clustering can be written as:
@@
We design a Semi-Definite Programming (SDP) relaxation and then round the SDP solution to a clustering. Let us denote the assignment variables as an $n\times k$ matrix $V$, such that each entry $V_{i,j}$ indicates whether a task $i$ belongs to a cluster $j$, for every $i=1,\dots,n$, $j=1,\dots,k$. Moreover, let the $j$th column of $V$, the characteristic vector of the $j$-th cluster, be denoted as $v_{j}$. Under this assignment, the sum of $V_{i,j}$ across any task $i$ must be one, as we allow one task to be assigned in a single group. By contrast, the sum of $V_{i,j}$ across $C_{j}$ is the number of tasks assigned to $C_{j}$, which is at least one.
@@
A natural question is using alternative algorithms such as spectral clustering or Lloyd’s clustering. We find that these algorithms are not as robust as the SDP relaxation because the scale of the loss values varies across rows for different tasks. We describe a toy example to illustrate. Suppose $T$ is a $6$ by $6$ matrix involving three clusters $C_{1},C_{2},C_{3}$ of size $2$ each. The affinity in $C_{1}$ is $7$, while the affinity scores in $C_{2}$ and $C_{3}$ are $20,19$, respectively. We find that both spectral clustering and Lloyd’s clustering will group $C_{2}$ and $C_{3}$ together, while the SDP relaxation manages to separate them. See Figure 2 for an illustration. For this reason, we use the SDP relaxation in Grad-TAG.
@@
A natural question is whether one can quantify the approximation ratio of the SDP relaxation (10). Although this is a well-studied problem in approximation algorithms (ailon2008aggregating), task affinity violates the metric condition typically required to obtain guarantees in this literature. In particular, the triangle inequality $T_{i,j}+T_{j,k}\geq T_{i,k}$ is violated. It is possible that by assuming intra-cluster separation (see, e.g., awasthi2015relax), one might be able to analyze the SDP theoretically. This is left for future work.
@@ 5. Experiments
Our experiments show that Grad-TAE approximates the true task affinities (based on full model training) within a relative error of less than 2.7%, using less than 3% of the computational cost of full training. Further, Grad-TAG achieves comparable downstream accuracy to existing methods in two canonical applications, multi-label classification on graphs and language model fine-tuning, while using 32.8$\times$ fewer FLOPs. Lastly, we discuss the parameters and the steps as part of our algorithm, including the comparison with alternative clustering.
@@ 5.1. Experimental Setup
We note that our algorithm applies to a wide range of multitask learning scenarios. For a representative evaluation, we focus on multi-label prediction on graphs and language model fine-tuning. In the first setting, each labeling task corresponds to a subgraph within a graph. Given a seed set of each labeling as the training set, the goal is to identify the remaining nodes of the subgraph. This can be cast as multitask learning by viewing each labeling as a binary classification task. The objective is to optimize the average accuracy of all the labeling tasks.
@@
For fine-tuning language models, we use two text classification datasets from SuperGLUE (wang2019superglue), specifically RTE and WiC. Each dataset includes 100 instructions, with ten sourced from bach2022promptsource and 90 generated using the automatic instruction generation method in (zhou2022large). Thus, each dataset has 100 tasks, each corresponding to fine-tuning with one instruction. We use T5-Base (roberts2019exploring) as the encoder for the MTL model. The choice of this encoder is without loss of generality, as we expect similar results to hold on other encoders.
@@
We assess the accuracy of estimated task affinity by measuring the distance between our estimated task affinities and those computed from fully trained models.
@@ 5.2. Task Affinity Estimation
We now report the results of our estimation procedure. We regard the task affinity scores computed from fully trained models as the target, denoted as ${T}^{\star}$. Then, after running Grad-TAE, we compute the affinity matrix $T$, and measure the relative distance between $T$ and ${T}^{\star}$ as:
@@
To explain our findings, we set the number of meta-initializations to $M=1$ and vary the projection dimension $d$ among 50, 100, 200, and 400. We note that all these values yield an estimation of $T^{\star}$ within 11% distance. As expected, increasing $d$ leads to better estimation. After $d$ increases above $200$, the distance metric stabilizes to around 5.7%. Thus, we set $d$ as 200 in the remaining experiments. As a remark, this is approximately $15\log(p)$, where $p=683,370$ in this experiment, aligning with our analysis in Proposition 3.3. Remarkably, under this setting, Grad-TAE uses 3.5 GPU hours and achieves 130$\times$ less computation compared to fully-trained models!
Next, we fix $d=200$ while increasing $M$ up to $9$. This further reduces the distance metric to 5.4%, with 45.0$\times$ less compute cost. We observe diminishing returns from the ensemble once $M$ goes beyond $5$. Thus, we will set $M$ as $5$ in the remaining experiments. This uses 17.6 GPU hours and 44.9$\times$ less computation than fully-trained models.
@@
We have shown that Grad-TAE significantly reduces the computational cost in task affinity estimation. To verify that these efficiency gains are consistent across different data modalities, we apply Grad-TAE to a text classification dataset (RTE) and an image classification dataset (DomainNet) (peng2019moment). The RTE data set contains 100 tasks. We use T5-Base and compute higher-order task affinity with 2000 subsets of size 10. The DomainNet data set contains six tasks. We use ResNet-50 and compute higher-order task affinity with 20 subsets of size 3. On the two data sets, our algorithm reduces computation by 42.6$\times$ and 9.5$\times$, respectively, compared to computing true higher-order task affinities while incurring less than 3% relative error. The smaller speedup in the image dataset is due to the fewer total models trained on task subsets.
@@ 5.3. Comparison for Task Grouping
We set up many baselines covering heuristic solutions and recent optimization techniques.
@@ 5.3.3. Fine-tuning language models
Next, we report the results from fine-tuning language models (T5 base) on text classification with $n=100$ instructions. Again, we use our algorithm to estimate higher-order task affinity scores and apply SDP clustering to group tasks. We illustrate our results in Figure 4(b) while deferring the complete comparison to Appendix C. We use $1-$ accuracy as the error rate on the text classification datasets. In particular, our algorithm outperforms single-task learning by 1.9%.
@@ 6. Conclusion
This paper designs an efficient estimation algorithm to compute task affinity scores. The main idea is to pre-train a meta-initialization on all tasks and then use the initialization’s gradients to estimate the fine-tuned model parameters for a particular task combination using logistic regression. A random projection is applied to the gradients to reduce the dimension of the regression. Then, we design a robust clustering algorithm to accompany the task affinity estimation, which together yields an efficient multitask learning algorithm. Experiments show that the algorithm can scale to as many as $500$ tasks on large graphs while accurately approximating the true task affinity scores. The overall algorithm gives the best tradeoff between computation and performance compared to existing multitask learning methods.
@@ Acknowledgement
Thanks to the anonymous referees for their comments. This research is partly supported by NSF award IIS-2412008 and Northeastern University’s Transforming Interdisciplinary Experiential Research (TIER) 1: Seed Grant/Proof of Concept Program. D. Li was also partially funded by JPMorgan Chase & Co. Any views or opinions expressed herein are solely those of the authors listed, and may differ from the views and opinions expressed by JPMorgan Chase & Co. or its Affiliates.